Lavage Rapide is a Canadian company that owns and operates a large automatic car wash facility near Montreal. The following table provides data concerning the company’s costs:

Fixed Cost
per Month
Cost per
Car Washed
Cleaning supplies $ 0.70
Electricity $ 1,400 $ 0.09
Maintenance $ 0.30
Wages and salaries $ 4,700 $ 0.40
Depreciation $ 8,200
Rent $ 1,900
Administrative expenses $ 1,700 $ 0.02

For example, electricity costs are $1,400 per month plus $0.09 per car washed. The company actually washed 8,400 cars in August and collected an average of $6.60 per car washed.

Required:

Prepare the company’s flexible budget for August.

Lavage Rapide is a Canadian company that owns and operates a large automatic car wash facility near Montreal. The following table provides data concerning the company’s costs:

Fixed Cost
per Month
Cost per
Car Washed
Cleaning supplies $ 0.70
Electricity $ 1,400 $ 0.09
Maintenance $ 0.30
Wages and salaries $ 4,700 $ 0.40
Depreciation $ 8,200
Rent $ 1,900
Administrative expenses $ 1,700 $ 0.02

For example, electricity costs are $1,400 per month plus $0.09 per car washed. The company actually washed 8,400 cars in August and collected an average of $6.60 per car washed.

Required:

Prepare the company’s flexible budget for August.

Lavage Rapide
Flexible Budget
For the Month Ended August 31
Revenue
Expenses:
Cleaning supplies
Electricity
Maintenance
Wages and salaries
Depreciation
Rent
Administrative expenses
Total expense 0
Net operating income $0
Answer #1

Lavage Rapide

Flexible Budget

For the Month Ended August 31

Revenue

$ 55,440.00

Expenses:

Cleaning supplies

$ 5,880.00

Electricity

$ 2,156.00

Maintenance

$ 2,520.00

Wages and salaries

$ 8,060.00

Depreciation

$ 8,200.00

Rent

$ 1,900.00

Administrative expenses

$ 1,868.00

Total expense

$ 30,584.00

Net operating income

$ 24,856.00

Working

Lavage Rapide

Flexible Budget

For the Month Ended August 31

Revenue

=8400 x 6.6

Expenses:

Cleaning supplies

=8400 x 0.7

Electricity

=1400+(8400 x 0.09)

Maintenance

=(8400 x 0.3)

Wages and salaries

=4700+(8400 x 0.4)

Depreciation

8200

Rent

1900

Administrative expenses

=1700+(8400 x 0.02)